# DWGS: Enhancing Sparse-View Gaussian Splatting with Hybrid-Loss Depth Estimation and Bidirectional Warping

### v1

@@ III-D3 Virtual View Selection & Fusion
$I^{\prime}_{vir}=Fuse_{k}(I_{vir}|Order(Score(R_{src},R_{vir}))).$ (21)
@@
Figure 6: Multi-View Fusion. Fusing the two left virtual views produces the result shown on the right.
@@ III-E Occlusion-Aware Reconstruction
In NVS, models are typically optimized to reconstruct only the visible regions present in the training views. However, synthesizing views beyond the original training set often reveals occluded areas, presenting a significant challenge. This issue is exacerbated under sparse-view conditions, where limited coverage increases the probability of exposing unseen regions in novel views.
@@
$M_{local\_FG}=\mathcal{F}(Soble(D_{mono}),Slice(D_{mono})).$ (24)
@@
$\mathcal{S}(t)=\begin{cases}Train(I_{inp},P_{BG}),&if\ t\leq t_{BG},\\ Train(I,\ P),&if\ t>t_{BG}\end{cases}$ (27)
@@ IV-A1 Datasets
DWGS is evaluated on five widely adopted public datasets: LLFF, IBRNet, Blender, DTU, and Tanks&Temples. These datasets cover forward-facing, synthetic, object-centric, and 360° unbounded scenes, ensuring a comprehensive assessment of robustness across diverse settings. For LLFF and IBRNet (forward-facing scenes), we follow the protocol of DNGaussian [27] and SCGaussian [37]: one every 8 images is held out for testing, while 3 training views are uniformly sampled from the remainder. Evaluation is performed over LLFF scenes with 8×, 4× downsampling rate and IBRNet scenes with 2× downsampling rate. For Blender (synthetic objects), models are trained on 8 views and evaluated on 25 uniformly sampled test images with a 2× downsampling rate. For DTU (object-centric scenes), we adopt the setup from SCGaussian [37], using provided object masks during evaluation to exclude background regions and focus on the target object. A 4× downsampling rate is applied. For Tanks&Temples (360° scenes), 6 scenes are selected. Training uses 24 views, and training and testing split strategy follows the same protocol as used in LLFF (one every 8 images) with 4× downsampling rate.
3DGS          SID          MCGS        SCGaussian        DWGS(ours)         GT
Figure 8: Qualitative comparison on LLFF dataset under 3 training views. The rendering results of our DWGS are more accurate and display finer details.
3DGS          FSGS        DNGaussian       SCGaussian       DWGS(ours)         GT
Figure 9: Qualitative comparison on IBRNet dataset with 3 training views. The rendering results of DWGS method are more accurate and display finer details.
@@ IV-A2 Evaluation Metrics
Rendering quality is evaluated using PSNR, SSIM, and LPIPS. In addition, following DNGaussian [27], the Average (AVG) is computed as a composite metric, defined as the geometric mean of MSE ($10^{-\text{PSNR}/10}$), $\sqrt{1-\text{SSIM}}$, and LPIPS. Efficiency is assessed via frames per second (FPS) and total training time. To ensure statistical robustness, each scene is evaluated 7 times. The highest and lowest PSNR values are discarded, and the remaining 5 results are averaged to obtain the scene-level performance. Final dataset-level metrics are reported as the average across all scenes.
@@ IV-A3 Baselines
DWGS is compared against several excellent sparse-view NVS jobs, including NeRF-based and 3DGS-based approaches. NeRF-based baselines include FreeNeRF[8] and SparseNeRF[7]. 3DGS-based baselines include vanilla 3DGS[5], FSGS[26], SID[39], DNGaussian [27], MCGS[36], and SCGaussian [37]. For NeRF-based methods, we report their best quantitative results from the respective papers. To ensure a fair comparison, we use public code with consistent training-testing splits and a unified base environment with core dependencies fixed, including PyTorch and CUDA.
@@ IV-A4 Implementation Details
DWGS is built upon the 3DGS framework[5] and implemented in PyTorch. Camera poses are estimated using COLMAP across all views for robustness. To ensure consistent comparison, similar to SCGaussian, we employed the GIM-DKM[42] model for dense matching to extract matching points per training view. Besides, we use ML-depth-pro [43] for monocular depth estimation, which excels in capturing fine details. Both depth estimation and scene training are iterated 2,000 times, following the configuration of SCGaussian [37]. Initialize and generate Gaussian primitives after filtering outliers. Bidirectional Warping generates 50 virtual views; one virtual view iteration is performed for every 5 real view iterations. The Occlusion-Aware Reconstruction applies a depth difference mask to prioritize rendering of occluded region Gaussians in the first 1,000 iterations, followed by training all Gaussians. Gaussian cloning and pruning follow the vanilla 3DGS settings. All experiments are conducted on a single NVIDIA RTX 3090 GPU. For detailed configuration, please refer to the publicly available code.
@@
FreeNeRF[8] 19.63 0.612 0.308 0.134 18.73 0.562 0.384 0.169 19.76 0.588 0.333 0.135
SparseNeRF[7] 19.86 0.624 0.328 0.127 19.07 0.564 0.401 0.168 19.90 0.593 0.364 0.137
3DGS[5] 15.42 0.383 0.463 0.218 13.28 0.350 0.486 0.264 17.20 0.556 0.355 0.165
FSGS[26] 20.27 0.697 0.206 0.102 19.70 0.667 0.265 0.117 19.67 0.605 0.306 0.127
SID[39] 20.40 0.701 0.215 0.102 19.08 0.635 0.336 0.135 19.44 0.599 0.360 0.137
DNGaussian[27] 19.12 0.591 0.294 0.132 18.23 0.575 0.386 0.155 18.14 0.554 0.415 0.161
MCGS[36] 20.32 0.700 0.219 0.103 19.63 0.663 0.292 0.122 21.02 0.674 0.282 0.109
SCGaussian[37] 20.73 0.725 0.196 0.095 20.03 0.683 0.266 0.114 21.47 0.689 0.275 0.103
DWGS(Ours) 21.13 0.735 0.189 0.090 20.30 0.693 0.256 0.109 22.19 0.708 0.249 0.093
@@ IV-B1 LLFF & IBRNet
Quantitative results on the LLFF and IBRNet datasets under extremely sparse settings (3 training views) are summarized in Table II. DWGS achieves state-of-the-art performance, reaching a PSNR of 21.13 dB and LPIPS of 0.189 on LLFF, and 22.19 dB PSNR with 0.249 LPIPS on IBRNet, outperforming all baseline methods across metrics.
@@
In contrast, DWGS produces renderings that closely match the GT across all forward-facing scenes, with enhanced edge sharpness, improved texture preservation, and minimal artifacts. These results demonstrate DWGS’s robustness and superiority in challenging sparse-view reconstruction of forward-facing scenes.
@@ IV-B2 Blender
Quantitative comparisons on the Blender dataset with 8 training views are presented in Table III, DWGS achieves the best performance across all metrics.
Qualitative results are shown in Figure 10. In the Chair object, MCGS produces noticeable floaters, while DNGaussian suffers from blurred details. For the Lego object, both MCGS and DNGaussian fail to reconstruct fine structures, leading to a loss of detail. In the Drums and Ship objects, 3DGS and SCGaussian exhibit floters, with 3DGS showing the most severe floters as well as a clear color discrepancy from the GT. In contrast, DWGS produces superior reconstruction results and more accurate depth maps, with sharper geometry and better-preserved textures.
MCGS     DNGaussian    DWGS(ours)      GT
3DGS     SCGaussian    DWGS(ours)      GT
Figure 10: Qualitative comparison on Blender dataset with 8 training views. DWGS method excels both in geometry and rendering qualities.
TABLE III: Quantitative comparisons on the Blender (1/2 resolution) datasets with 8 training views. Top-3 entries are highlighted: red (1st), orange (2nd), yellow (3rd).
@@
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ AVG $\downarrow$
3DGS[5] 22.85 0.836 0.141 0.066
FSGS[26] 16.35 0.601 0.369 0.175
SID[39] 12.65 0.729 0.321 0.208
DNGaussian[27] 22.79 0.870 0.106 0.058
MCGS[36] 24.06 0.887 0.089 0.048
SCGaussian[37] 23.33 0.883 0.097 0.053
DWGS(Ours) 24.18 0.894 0.086 0.047
3DGS      DNGaussian     DWGS(ours)      GT
FSGS       SID       DWGS(ours)      GT
MCGS     SCGaussian     DWGS(ours)      GT
@@ IV-B3 DTU
Quantitative comparisons on the DTU dataset are summarized in Table IV, where DWGS achieves the highest scores across all metrics, followed by SCGaussian and MCGS.
Figure 11 presents the rendering results. Both 3DGS and DNGaussian produce a considerable number of floaters. FSGS and SID fail to reconstruct textual details and suffer from overall blurriness. While MCGS and SCGaussian also exhibit floaters, MCGS shows the most severe cases, which further lead to inaccurate depth estimations. Benefiting from the outlier filtering mechanism and virtual view constraints, DWGS significantly reduces artifacts and improves texture fidelity, particularly around text and fine structures.
TABLE IV: Quantitative comparisons on the DTU (1/4 resolution) datasets with 3 training views. Top-3 entries are highlighted: red (1st), orange (2nd), yellow (3rd).
@@
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ AVG $\downarrow$
FreeNeRF[8] 19.92 0.787 0.182 0.098
SparseNeRF[7] 19.55 0.769 0.201 0.102
3DGS[5] 12.76 0.595 0.376 0.233
FSGS[26] 17.41 0.728 0.247 0.132
SID[26] 16.35 0.654 0.334 0.165
DNGaussian[27] 18.46 0.807 0.168 0.101
MCGS[36] 19.02 0.810 0.154 0.094
SCGaussian[37] 19.11 0.857 0.123 0.082
DWGS(Ours) 20.22 0.872 0.110 0.072
@@ IV-B4 Tanks&Temples
The Tanks&Temples dataset is employed to evaluate the performance of DWGS in large-scale 360° unbounded scenes. Quantitative results are presented in Table V. Note that evaluations on this dataset utilize only the Hybrid-Loss Depth Estimation component of our framework. DNGaussian and MCGS are excluded from comparison as they lack configurations suitable for 360° unbounded scenes.
Qualitative results are shown in Figure 12. As SID is an enhanced version of FSGS, only FSGS is presented for clarity. In the Family and Horse scenes, both vanilla 3DGS and SID exhibit significant artifacts, with the latter also showing noticeable color rendering errors. SCGaussian suffers from partial object omission. In contrast, DWGS produces more complete and geometrically stable reconstructions with reduced artifacts, demonstrating its robustness in challenging large-scale unbounded scenarios.
3DGS             SID           SCGaussian           DWGS(ours)          GT
Figure 12: Qualitative comparison on Tank&Temple dataset with 24 training views. DWGS method excels both in geometry and rendering qualities in difficult areas.
TABLE V: Quantitative comparisons on the Tank&Temple (1/4 resolution) datasets with 24 training views. Top-3 entries are highlighted: red (1st), orange (2nd), yellow (3rd).
Method Tank&Temple
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ AVG $\downarrow$
3DGS[5] 18.77 0.706 0.267 0.124
FSGS[26] 19.39 0.720 0.273 0.118
SID[26] 19.51 0.723 0.273 0.117
SCGaussian[37] 18.96 0.722 0.298 0.125
DWGS(Ours) 20.00 0.749 0.268 0.110

### v2

@@ III-D3 Virtual View Selection & Fusion
$I^{\prime}_{vir}=Fuse_{k}(I_{vir}|Order(Score(R_{src},R_{vir}))),$ (21)
@@
Figure 6: Multi-View Fusion. Fusing the left, middle virtual views produces the result shown on the right.
@@ III-E Occlusion-Aware Reconstruction
In NVS, models are typically optimized to reconstruct only the visible regions present in the training views. However, synthesizing views beyond the original training set often reveals occluded areas. This issue is exacerbated under sparse-view conditions, where limited coverage increases the probability of exposing unseen regions in novel views.
@@
$M_{local\_FG}=\mathcal{F}(Soble(D_{mono}),Slice(D_{mono})),$ (24)
@@
$\mathcal{S}(t)=\begin{cases}Train(I_{inp},P_{BG}),&if\ t\leq t_{BG},\\ Train(I,\ P),&if\ t>t_{BG},\end{cases}$ (27)
@@ IV-A1 Datasets
HBSplat is evaluated on five widely adopted public datasets: LLFF, IBRNet, Blender, DTU, and Tanks&Temples. These datasets cover forward-facing, synthetic, object-centric, and 360° unbounded scenes, ensuring a comprehensive assessment of robustness across diverse settings. For LLFF and IBRNet (forward-facing scenes), we follow the protocol of DNGaussian [27]: one every 8 images is held out for testing, while 3 training views are uniformly sampled from the remainder. Evaluation is performed over LLFF scenes with 8×, 4× downsampling rate and IBRNet scenes with 2× downsampling rate. For Blender (synthetic objects), models are trained on 8 views and evaluated on 25 uniformly sampled test images with 2× downsampling rate. For DTU (object-centric scenes), we adopt the setup from SCGaussian [37], using provided object masks during evaluation to exclude background regions and focus on the target object. 4× downsampling rate is applied. For Tanks&Temples (360° scenes), 6 scenes are selected. Training uses 24 views, training and testing split strategy follows the same protocol as used in LLFF (one every 8 images) with 4× downsampling rate.
3DGS          SID          MCGS        SCGaussian        HBSplat(Ours)         GT
Figure 8: Qualitative comparison on LLFF dataset under 3 training views. The rendering results of our HBSplat are more accurate and display finer details.
3DGS          FSGS        DNGaussian       SCGaussian       HBSplat(Ours)         GT
Figure 9: Qualitative comparison on IBRNet dataset with 3 training views. The rendering results of HBSplat method are more accurate and display finer details.
@@ IV-A2 Evaluation Metrics
Rendering quality is evaluated using PSNR, SSIM, and LPIPS. In addition, following DNGaussian [27], the Average (AVG) is computed as a composite metric, defined as the geometric mean of MSE ($10^{-\text{PSNR}/10}$), $\sqrt{1-\text{SSIM}}$, and LPIPS. Efficiency is assessed via frames per second (FPS) and average training time. Final dataset-level metrics are reported as the average across all scenes.
@@ IV-A3 Baselines
HBSplat is compared against several excellent sparse-view NVS jobs, including NeRF-based and 3DGS-based approaches. NeRF-based baselines include FreeNeRF[8] and SparseNeRF[7]. 3DGS-based baselines include vanilla 3DGS[5], FSGS[26], SID[39], DNGaussian [27], MCGS[36], and SCGaussian [37]. For NeRF-based methods, we report their best quantitative results from the respective papers. To ensure a fair comparison, we use public code with consistent training-testing splits and a unified base environment with core dependencies fixed, including PyTorch and CUDA.
@@ IV-A4 Implementation Details
HBSplat is built upon the 3DGS framework[5]. Camera poses are estimated using COLMAP across all views for robustness. To ensure consistent comparison, similar to SCGaussian, we employed the GIM-DKM[44] model for dense matching to extract matching points per training view. Besides, we use ML-depth-pro [45] for monocular depth estimation, which excels in capturing fine details. Both depth estimation and scene training are iterated 2,000 times. Initialize and generate Gaussian primitives after filtering outliers. Bidirectional Warping generates 50 virtual views; one virtual view iteration is performed for every 5 real view iterations. The Occlusion-Aware Reconstruction applies a depth difference mask to prioritize rendering of occluded region Gaussians in the first 1,000 iterations, followed by training all Gaussians. Gaussian cloning and pruning follow the vanilla 3DGS settings. All experiments are conducted on a single NVIDIA RTX 3090 GPU. Please refer to the publicly available code for detailed configuration.
@@
FreeNeRF
[CVPR23]
[8]
19.63 0.612 0.308 0.134 18.73 0.562 0.384 0.169 19.76 0.588 0.333 0.135
SparseNeRF
[ICCV23]
[7]
19.86 0.624 0.328 0.127 19.07 0.564 0.401 0.168 19.90 0.593 0.364 0.137
3DGS
[SIGGRAPH23]
[5]
15.42 0.383 0.463 0.218 13.28 0.350 0.486 0.264 17.20 0.556 0.355 0.165
FSGS
[ECCV24]
[26]
[…]
SID
[ICASSP25]
[39]
[…]
DNGaussian
[CVPR24]
[27]
19.12 0.591 0.294 0.132 18.23 0.575 0.386 0.155 18.14 0.554 0.415 0.161
MCGS
[TPAMI25]
[36]
[…]
SCGaussian
[NeurIPS24]
[37]
[…]
TABLE III: Quantitative comparisons on the NeRF-LLFF dataset with 2,3,5 training views. The best and second-best entries are marked in red and orange, respectively.
Scene Method 2-view 3-view 5-view
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ AVG $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ AVG $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ AVG $\downarrow$
Fern 3DGS
[SIGGRAPH23]
14.07 0.306 0.509 0.255 15.80 0.348 0.502 0.220 25.02 0.647 0.255 0.078
SCG
[NeurIPS24]
19.38 0.627 0.317 0.133 22.15 0.741 0.175 0.084 24.20 0.824 0.120 0.057
HBSplat(Ours) 20.09 0.658 0.258 0.115 22.73 0.753 0.173 0.078 24.52 0.832 0.110 0.054
Flower 3DGS
[SIGGRAPH23]
16.15 0.370 0.437 0.203 16.37 0.445 0.437 0.195 26.10 0.782 0.146 0.055
SCG
[NeurIPS24]
19.84 0.608 0.307 0.134 21.66 0.690 0.217 0.103 25.86 0.858 0.097 0.045
HBSplat(Ours) 20.10 0.630 0.274 0.126 21.70 0.693 0.221 0.104 25.96 0.861 0.090 0.044
Fortress 3DGS
[SIGGRAPH23]
15.70 0.325 0.379 0.203 18.80 0.437 0.367 0.153 23.74 0.613 0.229 0.084
SCG
[NeurIPS24]
21.34 0.558 0.294 0.114 25.00 0.823 0.122 0.056 27.48 0.867 0.094 0.040
HBSplat(Ours) 22.17 0.646 0.274 0.097 25.68 0.836 0.117 0.051 27.60 0.871 0.093 0.039
Horns 3DGS
[SIGGRAPH23]
14.35 0.291 0.511 0.250 15.65 0.391 0.463 0.214 19.07 0.528 0.359 0.145
SCG
[NeurIPS24]
16.84 0.554 0.372 0.176 19.86 0.737 0.221 0.109 22.71 0.821 0.158 0.077
HBSplat(Ours) 17.76 0.624 0.319 0.152 20.33 0.746 0.219 0.105 23.38 0.832 0.145 0.070
Leaves 3DGS
[SIGGRAPH23]
13.18 0.249 0.452 0.266 14.94 0.342 0.410 0.220 18.20 0.528 0.300 0.145
SCG
[NeurIPS24]
15.30 0.419 0.454 0.216 17.97 0.651 0.240 0.131 19.46 0.720 0.203 0.106
HBSplat(Ours) 15.86 0.482 0.396 0.192 18.14 0.659 0.237 0.128 19.70 0.734 0.194 0.102
Orchids 3DGS
[SIGGRAPH23]
14.88 0.214 0.492 0.242 15.95 0.313 0.460 0.213 20.21 0.472 0.340 0.132
SCG
[NeurIPS24]
14.40 0.384 0.363 0.218 16.43 0.530 0.248 0.157 17.93 0.619 0.203 0.126
HBSplat(Ours) 14.63 0.414 0.328 0.205 16.67 0.543 0.234 0.150 17.83 0.617 0.202 0.127
Room 3DGS
[SIGGRAPH23]
12.09 0.326 0.581 0.308 13.51 0.451 0.546 0.262 16.23 0.451 0.508 0.207
SCG
[NeurIPS24]
19.87 0.793 0.228 0.104 22.12 0.858 0.150 0.071 26.20 0.910 0.107 0.045
HBSplat(Ours) 20.75 0.823 0.190 0.089 22.53 0.868 0.141 0.066 26.64 0.915 0.103 0.043
Trex 3DGS
[SIGGRAPH23]
11.20 0.225 0.552 0.332 12.35 0.344 0.518 0.290 14.67 0.436 0.436 0.223
SCG
[NeurIPS24]
18.40 0.665 0.291 0.138 20.62 0.774 0.193 0.093 23.24 0.851 0.132 0.063
HBSplat(Ours) 19.35 0.702 0.236 0.115 21.27 0.787 0.171 0.084 24.01 0.864 0.120 0.057
Mean 3DGS
[SIGGRAPH23]
13.95 0.288 0.489 0.255 15.42 0.383 0.463 0.218 20.41 0.557 0.322 0.124
SCG
[NeurIPS24]
[…]
@@ IV-B1 LLFF & IBRNet
Quantitative results on the LLFF and IBRNet datasets under extremely sparse settings (3 training views) are summarized in Table II. HBSplat achieves state-of-the-art performance, reaching a PSNR of 21.13 dB and LPIPS of 0.189 on LLFF, and 22.19 dB PSNR with 0.249 LPIPS on IBRNet, outperforming all baseline methods across metrics. Detailed per-scene results on LLFF (III) further validate the advantage of HBSplat in forward-facing scene reconstruction.
@@
In contrast, HBSplat produces renderings that closely match the GT across all forward-facing scenes, with enhanced edge sharpness, improved texture preservation, and minimal artifacts. These results demonstrate HBSplat’s robustness and superiority in challenging sparse-view reconstruction of forward-facing scenes.
@@ IV-B2 Blender
Quantitative comparisons on the Blender dataset with 8 training views are presented in Table IV, where HBSplat achieves competitive results across all metrics.
MCGS     DNGaussian    HBSplat(Ours)      GT
3DGS     SCGaussian    HBSplat(Ours)      GT
Figure 10: Qualitative comparison on Blender dataset with 8 training views. HBSplat method excels both in geometry and rendering qualities.
TABLE IV: Quantitative comparisons on the Blender (1/2 resolution) datasets with 8 training views. Top-3 entries are highlighted: red (1st), orange (2nd), yellow (3rd).
@@
[…]
FreeNeRF
[CVPR23]
[8]
[…]
SparseNeRF
[ICCV23]
[7]
22.41 0.861 0.119 0.063
3DGS
[SIGGRAPH23]
[5]
22.85 0.836 0.141 0.066
FSGS
[ECCV24]
[26]
16.35 0.601 0.369 0.175
SID
[ICASSP25]
[39]
12.65 0.729 0.321 0.208
DNGaussian
[CVPR24]
[27]
22.79 0.870 0.106 0.058
MCGS
[TPAMI25]
[36]
[…]
SCGaussian
[NeurIPS24]
[37]
[…]
Qualitative results are shown in Figure 10. In the Chair object, MCGS produces noticeable floaters, while DNGaussian suffers from blurred details. For the Lego object, both MCGS and DNGaussian fail to reconstruct fine structures, leading to a loss of detail. In the Drums and Ship objects, 3DGS and SCGaussian exhibit floters, with 3DGS showing the most severe floters as well as a clear color discrepancy from the GT. In contrast, HBSplat produces superior reconstruction results and more accurate depth maps, with sharper geometry and better-preserved textures.
3DGS      DNGaussian     HBSplat(Ours)      GT
FSGS       SID       HBSplat(Ours)      GT
MCGS     SCGaussian     HBSplat(Ours)      GT
@@ IV-B3 DTU
Quantitative comparisons on the DTU dataset are summarized in Table V, where HBSplat achieves the highest scores across all metrics, followed by SCGaussian and MCGS.
Figure 11 presents the rendering results. Both 3DGS and DNGaussian produce a considerable number of floaters. FSGS and SID fail to reconstruct textual details and suffer from overall blurriness. While MCGS and SCGaussian also exhibit floaters, MCGS shows the most severe cases, which further lead to inaccurate depth estimations. Benefiting from the Outlier Filtering Mechanism and virtual view constraints, HBSplat significantly reduces artifacts and improves texture fidelity, particularly around text and fine structures.
TABLE V: Quantitative comparisons on the DTU (1/4 resolution) datasets with 3 training views. Top-3 entries are highlighted: red (1st), orange (2nd), yellow (3rd).
@@
[…]
FreeNeRF
[CVPR23]
[8]
19.92 0.787 0.182 0.098
SparseNeRF
[ICCV23]
[7]
19.55 0.769 0.201 0.102
3DGS
[SIGGRAPH23]
[5]
12.76 0.595 0.376 0.233
FSGS
[ECCV24]
[26]
17.41 0.728 0.247 0.132
SID
[ICASSP25]
[26]
16.35 0.654 0.334 0.165
DNGaussian
[CVPR24]
[27]
18.46 0.807 0.168 0.101
MCGS
[TPAMI25]
[36]
[…]
SCGaussian
[NeurIPS24]
[37]
[…]
@@ IV-B4 Tanks&Temples
The Tanks&Temples dataset is employed to evaluate the performance of HBSplat in large-scale 360° unbounded scenes. Quantitative results are presented in Table VI. Note that evaluations on this dataset utilize only the Hybrid-Loss Depth Estimation component of our framework. DNGaussian and MCGS are excluded from comparison as they lack configurations suitable for 360° unbounded scenes.
Qualitative results are shown in Figure 12. As SID is an enhanced version of FSGS, only SID is presented for clarity. In the Family and Horse scenes, both vanilla 3DGS and SID exhibit significant artifacts, with the latter also showing noticeable color rendering errors. SCGaussian suffers from partial object omission. In contrast, HBSplat produces more complete and geometrically stable reconstructions with reduced artifacts, demonstrating its robustness in challenging large-scale unbounded scenarios.
3DGS             SID            SCGaussian          HBSplat(Ours)          GT
Figure 12: Qualitative comparison on Tanks&Temples dataset with 24 training views. HBSplat method excels both in geometry and rendering qualities.
TABLE VI: Quantitative comparisons on the Tanks&Temples (1/4 resolution) datasets with 24 training views. Top-3 entries are highlighted: red (1st), orange (2nd), yellow (3rd).
Method Tanks&Temples
[…]
3DGS
[SIGGRAPH23]
[5]
18.77 0.706 0.267 0.124
FSGS
[ECCV24]
[26]
[…]
SID
[ICASSP25]
[26]
[…]
SCGaussian
[NeurIPS24]
[37]
[…]